An unbalanced four wire star connected load has a balanced voltage of 400v. The load are Z1= 4+j8, Z2= 3+j4, Z3= 15+j20 ohms.


A.) Find the line current

B.) Find the neutral current


1
Expert's answer
2021-01-18T01:17:06-0500

A) Find the line currents:


"I=\frac VZ,\\\space\\
I_1=\frac{V_1}{Z_1}=\frac{400/\sqrt3}{4+j8}=(11.5-j23.1)\text{ A}.\\\space\\

I_2=\frac{V_2}{Z_2}=\frac{-115+j200}{3+j4}=(18.1+j42.5)\text{ A}.\\\space\\
I_3=\frac{V_3}{Z_3}=\frac{-115-j200}{15+j20}=(-9.17-j1.10)\text{ A}."

B) Find the neutral current:


"I_N=I_1+I_2+I_3=(20.4+j18.3)\text{ A}."